Find the value of the point charge q0 if it has a voltage of 5.13 V and electric potential energy of 4.35 x 10-5 J


1
Expert's answer
2022-05-23T10:18:30-0400
"E_p=q_0V"

"q_0=\frac{E_p}{V}\\
=\frac{4.35*10^{-5}\:\rm J}{5.13\:\rm V}=8.48*10^{-6}\:\rm C"